OpenStudy (anonymous):

At a conference for 500 people, 10% of the participants are French, 35% are Americans, 20% are Germans, and are of other nationalities. How many French participants are at the conference?

OpenStudy (anonymous):

metal and fan for answer

OpenStudy (anonymous):

If there are 10% French, that means that \[\large \frac{1}{10}\] of everyone is French. So what we're gonna do is multiply our fraction, 1/10 by 500. So \[\large \frac{1}{10}*500\] Do you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(1/10)*500=50, correct?

OpenStudy (anonymous):

Yes, that would be it. So 50 would be your answer :)

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]